# Unraveling Instance Associations: A Closer Look for Audio-Visual Segmentation

### v5

@@ 1 Introduction
A particularly interesting AVC task is the audio-visual segmentation (AVS) [63, 64] that aims to segment all pixels of the sounding visual objects using a fully supervised model. A major challenge in AVS is achieving cross-modal alignment between sound and visual objects [50]. Current datasets poorly establish and evaluate this alignment, leading to undesired system behaviour and less effective evaluation. For instance, the dataset in [63] shows a “commonsense” bias because it assumes that certain objects are always the sound source in some scenarios. Fig. 1 shows an example of a scene from [63] with a bias toward the segmentation of the helicopter, even though other sound sources (e.g., person’s speech or bird’s singing) are plausible. Such biases can reduce cross-modal alignment understanding. Another challenge in AVS datasets [63] is the localization of the sound-producing object when multiple instances of the same object class are present. While visual cues (e.g., motion or visual semantics) can help, some actions with silent audio can introduce false positives [40]. Spatial audio can also be helpful [58, 47], as shown in Fig. 3, where we are more likely to segment the dog on the right if the spatial audio suggests that the sounding object is located on the right-hand side of the image. Regrettably, addressing the AVS dataset challenges mentioned above by collecting new datasets can be exorbitantly expensive. Therefore, we consider alternative data collection procedures to mitigate the problems described above and to effectively enable the training and evaluation of AVS methods that can better generalise to diverse AVS conditions. Ziqin (Quinn) Zhou
@@ 4.1 Preliminaries about Cross-Attention
Table 1: Quantitative (mIoU, $F_{\beta})$ audio-visual segmentation results (in %) on AVSBench dataset [63, 64] (resized to 224$\times$224) with ResNet50 [19] backbone. Best results in bold, second best underlined. Improvements against the second best are in the brackets.
@@
  Transformer   CATR [26]   74.80 86.60 52.80 65.30 - -
AuTR [31]   75.00 85.20 49.40 61.20 - -
AQFormer [24]   77.00 86.40 55.70 66.90 - -
AVSegFormer [16]   76.54 84.80 49.53 62.80 24.93 29.30
AVSC [29]   77.02 85.24 49.58 61.51 - -
BAVS [30]   77.96 85.29 50.23 62.37 24.68 29.63
  Per-pixel Classification   TPAVI [63]   72.79 84.80 47.88 57.80 20.18 25.20
AVSBG [18]   74.13 85.40 44.95 56.80 - -
ECMVAE [36]   76.33 86.50 48.69 60.70 - -
DiffusionAVS [35]   75.80 86.90 49.77 62.10 - -
Ours   85.77 (+7.81) 92.86 (+5.96) 62.39 (+6.69) 73.62 (+6.72) 44.70 (+19.77) 57.76 (+28.13)
@@ 4.2 Contrastive Audio-Visual Pairing (CAVP)
The ultimate goal of our contrastive learning is to discriminate the positive fusion of feature representations from the same semantic class and the negative fusion of feature representations from different semantic classes. Previous works [25, 34] suggest that the discrimination between samples is critical for contrastive learning, so simply drawing random negative samples from original audio-visual pairs could limit the representation learning effectiveness. To create an informative contrastive dataset, a practical way is to mine all possible combinations of audio-visual pairs in the latent space. However, it is not possible to create some of the negative pairs because we do not have access to all instance segmentation masks for each image (e.g., we cannot create $f_{CA}(\mathbf{u}_{v}^{cat},\mathbf{u}_{a}^{dog})$ in Fig. 4 because we do not have pixels labelled as “cat”). Motivated by our VPO procedure in Sec. 3, we observe that this barrier can be overcome by randomly paring audio and visual features within the batch, as shown in Fig. 4, offering us the potential to mine most of the conceivable audio-visual combinations.
@@
mIoU $\uparrow$
$F_{\beta}$uparrow
mIoU $\uparrow$
$F_{\beta}$uparrow
  TPAVI [63]
42.74
58.11
52.83
70.39
  $f_{CA}(\cdot)$
47.18
61.74
58.50
73.19
  SupCon
48.72
61.85
59.35
74.20
  CAVP
50.75
62.31
64.31
64.13
@@ 5.2 Results
Performance on resized AVSBench. We divide the performance comparison on AVSBench into the AVSBench-Objects (SS & MS) [63] and AVSBench-Semantics [64]. We compare the performance of SOTA methods with our CAVP in Tab. 1 using mIoU and $F_{\beta}$, which shows that our model surpasses the second-best methods w.r.t. mIoU by of $7.81\%$ on AVSBench-Object (SS) [63], $6.69\%$ on AVSBench-Object (MS) [63] and $19.77\%$ on AVSBench-Semantics [64] using ResNet50 [19] backbone. Please refer to the Supplementary Material for results obtained with the PVT-V2-B5 [57] backbone.
Performance on original resolution AVSBench. We introduce a new benchmark based on the AVSBench-Semantics dataset [64] in Tab. 2 to address the absence of the original resolution images in previous benchmarks shown in Tab. 1. In this new benchmark, we use AVSegFormer [16] [^2] https://github.com/vvvb-github/AVSegFormer and TPAVI [63] [^3] https://github.com/OpenNLPLab/AVSBench as the baseline methods [^4] We cannot include other methods from Tab. 1 as they were not publicly available online at the time of submission.. To save training and evaluation time, we train on the entire training set of AVSBench-Semantics and test the model on AVSBench-Semantics (SS) and AVSBench-Semantics (MS) subsets, as well as the entire testing set for AVSBench-Semantics to show partitioned model performance. The results on the entire AVSBench-Semantics in the last columns of Tables 1 and 2 show a significant mIoU improvement of $+16.55\%$ for AVSegFromer [16], $+23.21\%$ for TPAVI, and +$6.05\%$ for our method when compared to the results obtained with low image resolution under ResNet-50 backbone [19]. These results suggest the importance of using the original resolution in the AVSBench-Semantics dataset [64]. Also in Tab. 2, results show that our method consistently outperforms the baselines by a minimum of $4.13\%$ and $2.93\%$ improvements in mIoU and $F_{\beta}$, respectively. We show a visualisation of a 6-second video clip in Fig. 5 that displays a qualitative comparison between TPAVI, AVSegFormer and our CAVP. Notice how our method approximates the ground truth segmentation of the sounding objects more consistently than the other methods. Please refer to the Supplementary Material for more qualitative results.
@@ B.1 Implementation Details
Appendix C Attention Map Visualisation
C.1 T-SNE Visualisation
Table 7: Quantitative (mIoU, $F_{\beta})$ audio-visual segmentation results (in %) on AVSBench dataset [63, 64] (resized to 224$\times$224) with PVT-V2-B5 [57] backbone. The best results are in bold and the second best are underlined.
  PVT-V2-B5 [57]   Method   AVSBench-Object (SS) AVSBench-Object (MS) AVSBench-Semantics
  Transformer   CATR [26]   81.40 89.60 59.00 70.00 32.80 38.50
AuTR [31]   80.40 89.10 56.20 67.20 - -
AQFormer [24]   81.60 89.40 61.10 72.10 - -
AVSegFormer [16]   82.06 89.90 58.36 69.30 36.66 42.00
AVSC [29]   80.57 88.19 58.22 65.10 - -
BAVS [30]   81.96 88.60 58.63 65.49 32.64 36.42
  Per-pixel Classification   TPAVI [63]   78.74 87.90 54.00 64.50 29.77 35.20
AVSBG [18]   81.71 90.40 55.10 66.80 - -
ECMVAE [36]   81.74 90.10 57.84 70.80 - -
DiffusionAVS [35]   81.38 90.20 58.18 70.90 - -
Ours   87.33 (+5.27) 93.61 (+3.41) 67.31 (+6.21) 78.09 (+5.99) 48.59 (+11.93) 61.97 (+19.97)
@@ C.1 T-SNE Visualisation
C.2 Additional Results on Resized AVSBench
We follow previous methods [26, 31, 24, 16, 29, 30, 63, 18, 36, 35] to evaluate our model with PVT-V2-B5 [57] backbone on AVSBench-Objects (SS & MS) [63] and AVSBench-Semantics [64] with resized image resolution (224$\times$224). The results in Tab. 7 show that we improve mIoU by $5.27\%$, $6.21\%$ and $11.93\%$ on the respective three benchmarks.
C.3 Additional Results

### v7

@@ 1 Introduction
A particularly interesting AVC task is the audio-visual segmentation (AVS) [63, 64] that aims to segment all pixels of the sounding visual objects using a fully supervised model. A major challenge in AVS is achieving cross-modal alignment between sound and visual objects [50]. Current datasets poorly establish and evaluate this alignment, leading to undesired system behaviour and less effective evaluation. For instance, the dataset in [63] shows a “commonsense” bias because it assumes that certain objects are always the sound source in some scenarios. Fig. 1 shows an example of a scene from [63] with a bias toward the segmentation of the helicopter, even though other sound sources (e.g., person’s speech or bird’s singing) are plausible. Such biases can reduce cross-modal alignment understanding. Another challenge in AVS datasets [63] is the localisation of the sound-producing object when multiple instances of the same object class are present. While visual cues (e.g., motion or visual semantics) can help, some actions with silent audio can introduce false positives [40]. Spatial audio can also be helpful [58, 47], as shown in Fig. 3, where we are more likely to segment the dog on the right if the spatial audio suggests that the sounding object is located on the right-hand side of the image. Regrettably, addressing the AVS dataset challenges mentioned above by collecting new datasets can be exorbitantly expensive. Therefore, we consider alternative data collection procedures to mitigate the problems described above and to effectively enable the training and evaluation of AVS methods that can better generalise to diverse AVS conditions.
@@ 4.1 Preliminaries about Cross-Attention
Table 1: Quantitative ($\mathcal{J}$, $\mathcal{F}$) audio-visual segmentation results (in %) on AVSBench dataset [63, 64] (resized to 224$\times$224) with ResNet50 [19] backbone. Best results in bold, second best underlined. Improvements against the second best are in the brackets.
    ImageNet Pretrained Backbone
@@
            $\mathcal{J}$$&\mathcal{F}$uparrow $\mathcal{J}$uparrow $\mathcal{F}$uparrow $\mathcal{J}$$&\mathcal{F}$uparrow $\mathcal{J}$uparrow $\mathcal{F}$uparrow $\mathcal{J}$$&\mathcal{F}$uparrow $\mathcal{J}$uparrow $\mathcal{F}$uparrow
    Transformer     CATR [26]     80.70 74.80 86.60 59.05 52.80 65.30 - - -
        AuTR [31]     80.10 75.00 85.20 55.30 49.40 61.20 - - -
        AVSegFormer [16]     80.67 76.54 84.80 56.17 49.53 62.80 27.12 24.93 29.30
        AVSC [29]     81.13 77.02 85.24 55.55 49.58 61.51 - - -
        BAVS [30]     81.63 77.96 85.29 56.30 50.23 62.37 27.16 24.68 29.63
    Per-pixel Classification     TPAVI [63]     78.80 72.79 84.80 52.84 47.88 57.80 22.69 20.18 25.20
        AVSBG [18]     79.77 74.13 85.40 50.88 44.95 56.80 - - -
        ECMVAE [36]     81.42 76.33 86.50 54.70 48.69 60.70 - - -
        DiffusionAVS [35]     81.35 75.80 86.90 55.94 49.77 62.10 - - -
        Ours     83.84 78.78 88.89 61.48 55.82 67.14 32.83 30.37 35.29
    COCO Pretrained Backbone
    Transformer     AQFormer     81.70 77.00 86.40 61.30 55.70 66.90 - - -
    Per-pixel Cls.     Ours     83.75 78.72 88.77 62.34 56.42 68.25 - - -
@@ 4.2 Contrastive Audio-Visual Pairing (CAVP)
The ultimate goal of our contrastive learning is to discriminate the positive fusion of feature representations from the same semantic class and the negative fusion of feature representations from different semantic classes. Previous works [25, 34] suggest that the discrimination between samples is critical for contrastive learning, so simply drawing random negative samples from original audio-visual pairs could limit the representation learning effectiveness. A practical way to create an informative contrastive dataset is to mine all possible combinations of audio-visual pairs in the latent space. However, it is not possible to create some of the negative pairs because we do not have access to all instance segmentation masks for each image (e.g., we cannot create $f_{CA}(\mathbf{u}_{v}^{cat},\mathbf{u}_{a}^{dog})$ in Fig. 4 because we do not have pixels labelled as “cat”). Motivated by our VPO procedure in Sec. 3, we observe that this barrier can be overcome by randomly paring audio and visual features within the batch, as shown in Fig. 4, offering us the potential to mine most of the conceivable audio-visual combinations.
@@
Table 2: Quantitative (mIoU, $F_{\beta})$ audio-visual segmentation results (in %) on AVSBench dataset [63, 64] (resized to 224$\times$224) with ResNet50 [19] backbone. * indicate the our initial results.
    Per-Pixel Classification     CAVP*     85.77 92.86 62.39 73.62 44.70 57.76
        CAVP     89.43 94.50 72.79 83.05 44.70 57.76
@@
        mIoU $\uparrow$ $F_{\beta}$uparrow mIoU $\uparrow$ $F_{\beta}$uparrow
    TPAVI [63]     42.74 58.11 52.83 70.39
    $f_{CA}(\cdot)$     47.18 61.74 58.50 73.19
    SupCon     48.72 61.85 59.35 74.20
    CAVP     50.75 62.31 64.31 64.13
@@ 5.2 Results
Performance on resized AVSBench. We divide the performance comparison on AVSBench into the AVSBench-Objects (SS & MS) [63] and AVSBench-Semantics [64]. We compare the performance of SOTA methods with our CAVP in Tab. 1 using Jaccrd index ($\mathcal{J}$) and F-Score ($\mathcal{F}$) based on [63], which shows that our model surpasses the second-best methods w.r.t. mIoU by of $2.21\%$ on AVSBench-Object (SS) [63], $2.43\%$ on AVSBench-Object (MS) [63] and $5.68\%$ on AVSBench-Semantics [64] using ResNet50 [19] backbone. We separate the AQFormer [24] in the bottom of Tab. 1 as their backbone model utilizes pre-trained weights based on COCO [28] instead of ImageNet [13]. Additionally, we report model performance under conventional semantic segmentation evaluation metrics, same as Pascal VOC [15] and CityScape [12] as this method do not skip the no sounding frames during evaluation.
Performance on original resolution AVSBench. We introduce a new benchmark based on the AVSBench-Semantics dataset [64] in Tab. 3 to address the absence of the original resolution images in previous benchmarks shown in Tab. 1. In this new benchmark, we use AVSegFormer [16] [^2] https://github.com/vvvb-github/AVSegFormer and TPAVI [63] [^3] https://github.com/OpenNLPLab/AVSBench as the baseline methods [^4] We cannot include other methods from Tab. 1 as they were not publicly available online at the time of submission.. To save training and evaluation time, we train on the entire training set of AVSBench-Semantics and test the model on AVSBench-Semantics (SS) and AVSBench-Semantics (MS) subsets, as well as the entire testing set for AVSBench-Semantics to show partitioned model performance. The results on the entire AVSBench-Semantics in the last columns of Tables 1 and 3 show a significant mIoU improvement of $+16.55\%$ for AVSegFromer [16], $+23.21\%$ for TPAVI, and +$6.05\%$ for our method when compared to the results obtained with low image resolution under ResNet-50 backbone [19]. These results suggest the importance of using the original resolution in the AVSBench-Semantics dataset [64]. Also in Tab. 3, results show that our method consistently outperforms the baselines by a minimum of $4.13\%$ and $2.93\%$ improvements in mIoU and $F_{\beta}$, respectively. We show a visualisation of a 6-second video clip in Fig. 5 that displays a qualitative comparison between TPAVI, AVSegFormer and our CAVP. Notice how our method approximates the ground truth segmentation of the sounding objects more consistently than the other methods. Please refer to the Supplementary Material for more qualitative results.
@@ A.3 Strengths and Weaknesses of VPO
Appendix B Experiments
B.1 Implementation Details
@@
Appendix C Attention Map Visualisation
C.1 T-SNE Visualisation
C.2 Additional Results